### v6

# Structured Packing in LLM Training Improves Long Context Utilization

Abstract
Recent developments in long-context large language models have attracted considerable attention. Yet, their real-world applications are often hindered by ineffective context information use. This work shows that structuring training data to increase semantic interdependence is an effective strategy for optimizing context utilization. To this end, we introduce Structured Packing for Long Context (SPLiCe), a method for creating training examples by using information retrieval methods to collate mutually relevant documents into a single training context. We empirically validate SPLiCe on large $3$B and $7$B models, showing perplexity improvements and better long-context utilization on downstream tasks. Remarkably, already relatively short fine-tuning with SPLiCe is enough to attain these benefits. Additionally, the comprehensive study of SPLiCe reveals intriguing transfer effects such as training on code data leading to perplexity improvements on text data.
Machine Learning, Language Models, ICML
1 Introduction
Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI. Recently, the field has observed the rise of Long-Context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a). However, their ability to process long context is not always as effective as one hopes for. Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a). Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a). Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components, such as the architecture and training objectives, unchanged. The broad question is how to organize training data to enhance long context capabilities? Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved. The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization. We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25 or repository structure) to collate mutually relevant documents into a single training context.
Figure 1: Training examples generated by the standard sampling procedure vs proposed by SPLiCe. Similar colors indicate related documents, which could be found using a retrieval method (e.g., BM25 or contriver) or metadata (e.g., repository structure for code).
We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2 and $7$Bv2 (Geng & Liu, 2023) for only $2$B–$6$B tokens already brings perplexity reduction. This reduction translates to substantial improvements in handling long-context information in downstream tasks that require retrieval and in-context learning. These tasks include TREC (Li & Roth, 2002; Hovy et al., 2001), DBpedia (Lehmann et al., 2015), Qasper (Shaham et al., 2022; Dasigi et al., 2021), HotPotQA (Yang et al., 2018), and the lost-in-the-middle benchmark (Liu et al., 2023). We perform a comprehensive study of the design choices and properties of SPLiCe, showing that the acquired long-context capabilities transfer between modalities, such as code and text.
Our contributions can be summarized as follows:
  • We comprehensively show that structuring training data is a viable way of improving the long context utilization of LLMs. To this end, we introduce SPLiCe, a method for creating training examples by using retrieval to collate mutually relevant documents into a single training context.
  • We fine-tune OpenLLaMA $3$Bv2 and OpenLLaMA $7$Bv2 (Geng & Liu, 2023) using SPLiCe, showing that it improves perplexity on long context evaluation, in-context learning ability, and ability to retrieve relevant data from the context.
  • We present a thorough analysis of the design choices behind SPLiCe, such as the number of retrieved documents and the order in which documents are merged into a training example.
2 Method
SPLiCe is a method for constructing training examples that enhances the LLMs long-context utilization. This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long context language modeling. We make the source code available to ensure the reproducibility of our results. (For this anonymous submission we share the zip file with our code.)
Algorithm 1 SPLiCe training example construction
  Input:    $D$: document corpus   $k$: breadth hyper-parameter   $L$: maximum length of returned training example   RETRIEVE: retrieval method to use, e.g., BM25   ORDER: ordering method, e.g., id, or random shuffle
  Output: training example consisting of concatenated documents
  $d_{r}\sim D$Sample the root document}
  $D=D\setminus\{d_{r}\}$
  $C=[d_{r}]$
  $Q=\texttt{empty queue }$
  $Q.\texttt{PUSH}(d_{r})$
  while $Q\neq\emptyset$ and $\texttt{len}(C)\leq L$ do
     $d=Q.\texttt{POP}()$
     $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$
     {Retrieve top-$k$ most similar documents to $d$ using a selected method, e.g., BM25}
     for each $d_{i}$ in $d_{1},\ldots,d_{k}$ do
        if $d_{i}\in D$ then
           {RETRIEVE uses a precomputed index, and may return documents that are already in $C$}
           $C=C.\texttt{APPEND}(d_{i})$Append $d_{i}$ to $C$}
           $Q.\texttt{PUSH}(d_{i})$
           $D=D\setminus\{d_{i}\}$
        end if
     end for
  end while
  return $\texttt{TRIM}(\texttt{ORDER}(C),L)$
Rationale and intuitions
Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022). However, it is an open question how to achieve such benefits in pre-training or fine-tuning. The primary difficulty comes from the fact that long-range dependencies are rare in training data (de Vries, 2023) and diminish with distance. It is thus unlikely that a model will learn to utilize long context without more explicit guidance.
Recent studies indicate that structuring data, i.e., going beyond the i.i.d. paradigm, might be beneficial or even necessary to achieve good long-context performance. (Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example. (Chan et al., 2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance. Finally, (Shi et al., 2023b) shows that training on structured data improves the performance of LLMs on long-context tasks.
SPLiCe follows these intuitions. Namely, it constructs training examples by collating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context. We provide additional insights into the distributional properties of SPLiCe created data at the end of this section and in Appendix K.
Structured Packing for Long Context (SPLiCe)
We propose SPLiCe, a method that creates training examples by building a tree of documents, see Algorithm 1 and Figure 1 for a general overview.
SPLiCe starts by picking a random document from the dataset to create a single-node tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus. The final sequence consists of the tree flattened according to a certain tree traversal strategy.
The hyperparameter $k$ introduces flexibility and allows to interpolate between modes of retrieval. Indeed, for $k=1$ SPLiCe creates a path of related examples simulating a long document. For larger values of $k$, SPLiCe produces examples that are similar to the ones used by retrieval augmented models, e.g., (Borgeaud et al., 2022). Note that the dataset must be reasonably curated for SPLiCe to work efficiently. For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
Many possible retrieval methods can be used with SPLiCe (Retrieve function in Algorithm 1). In our experiments, we test the following:
  • SPLiCeRepo: based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together. A similar method has been pioneered by (Wu et al., 2022) and proposed in (Shi et al., 2023b) as an interesting future direction.
  • SPLiCeBM25: based on BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
  • SPLiCeCont: based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer model to rank documents based on their similarity to a query.
SPLiCe computational efficiency Given the dataset sizes used in training LLMs, the computational efficiency plays a non-trivial role. SPLiCeRepo is the fastest and easy to implement but requires additional directory structure, i.e., is not applicable to general web data. SPLiCeBM25 requires quadratic time index and needs dataset chunking to be applicable to large datasets, see Section 3.3. SPLiCeCont requires calculating embeddings for each document and retreval based on the cosine similarity. The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
Baseline The standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context. This is known as example packing (Brown et al., 2020).
Burstiness SPLiCe conjecturally falls into the framework presented in (Chan et al., 2022). This paper shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of ”burstiness”, i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence . In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the baseline.
3 Experiments with Medium-Scale Models
In this section, we conduct a comprehensive examination of the impact of document packing on long-context performance, using medium-scale models (see Section 4 for large models results). We demonstrate that structuring data with SPLiCe is a generally applicable approach to improve long context performance. We study the performance in Section 3.2 and design choices of SPLiCe in Section 3.3.
3.1 Experimental Setup
In this part, we employ $270$M models and the following protocol. Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023). Subsequently, we continue training using $1$B tokens with an extended context length of $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline. We employ the Focused Transformer (FoT) objective (Tworkowski et al., 2023) for context extension (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023). Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. This aims to prevent the model from overfitting to artificially created long documents.
In the evaluation, we measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by their inclusion of documents that can benefit from long-context information as demonstrated in (Chen et al., 2023) and (Li et al., 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length.
We refer to Appendix A for a detailed description of the model architecture. For information regarding the training and evaluation data, see Appendix G.
3.2 Experimental Results
In this section, we compare SPLiCe against Baseline. We test two retrieval approaches for SPLiCe: BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), and Contriever-MSMARCO (Izacard et al., 2022)) denoted as SPLiCeBM25 and SPLiCeCont respectively. In SPLiCeBM25 retrieval takes into account the whole content of the document, whereas in SPLiCeCont we use the first $512$ tokens and Faiss (Johnson et al., 2017) for fast approximate inner-product search (see Appendix H for details). By SPLiCeRepo, we denote the approach using the directory structure of code repositories. We also test our method with $64$K context length and report the results in Table 19 in Appendix J. In those experiments, we use $k=1$; see Section 3.3 for more details.
The results are presented in Table 3, Table 3, and Table 3, from which we draw the following conclusions.
Structuring data improves performance. We observe (see Table 3) that all variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset. This indicates that the structure of the data is important for long-context training.
The positive effects transfer between domains. We observe (see Table 3) positive effects when training on code and evaluating on arXiv and when training on web data and evaluating on code. This indicates that the positive effects of SPLiCe transfer between domains. We conjecture that better data mixtures could bring further benefits.
Retrieval method is of secondary importance. We observe (see Table 3) that the choice of retrieval method has a minor impact on the results. Typically, SPLiCeBM25 is slightly better than SPLiCeCont. On the code data, SPLiCeRepo is on par with SPLiCeBM25. We recall that SPLiCeRepo can be applied only to code data, whereas SPLiCeBM25 and SPLiCeCont are more general. In practice, we often observe that SPLiCeCont is faster than SPLiCeBM25 due to the use of Faiss.
[…]
Long Context Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
C SPLiCeBM25 5.46 (.09) 3.20 (.17) 2.81 (.12) 2.22 (.11) 2.94 (.13) 3.10 (.13)
SPLiCeCont 5.48 (.07) 3.22 (.15) 2.82 (.11) 2.23 (.10) 2.96 (.11) 3.11 (.12)
SPLiCeRepo 5.47 (.08) 3.22 (.15) 2.83 (.10) 2.24 (.09) 2.96 (.11) 3.12 (.11)
Baseline 5.55 3.37 2.93 2.33 3.07 3.23
C# SPLiCeBM25 5.52 (.13) 3.33 (.25) 2.90 (.17) 2.46 (.19) 3.11 (.20) 3.26 (.20)
SPLiCeCont 5.53(.12) 3.35(.23) 2.91(.16) 2.48(.17) 3.12(.19) 3.27(.19)
SPLiCeRepo 5.53 (.12) 3.35 (.23) 2.91 (.16) 2.49 (.16) 3.12 (.19) 3.27 (.19)
Baseline 5.65 3.58 3.07 2.65 3.31 3.46
Python SPLiCeBM25 5.47 (.10) 3.25 (.21) 2.53 (.09) 2.41 (.15) 3.02 (.15) 3.17 (.15)
SPLiCeCont 5.49 (.08) 3.28 (.18) 2.53 (.09) 2.43 (.13) 3.03 (.14) 3.19 (.13)
SPLiCeRepo 5.48 (.09) 3.27 (.19) 2.54 (.08) 2.44 (.12) 3.03 (.14) 3.18 (.14)
Baseline 5.57 3.46 2.62 2.56 3.17 3.32
Wikipedia SPLiCeBM25 5.64 (.09) 3.82 (.15) 3.26 (.11) 2.87 (.13) 3.55 (.13) 3.68 (.13)
SPLiCeCont 5.65 (.08) 3.87 (.10) 3.30 (.07) 2.92 (.08) 3.59 (.09) 3.72 (.09)
Baseline 5.73 3.97 3.37 3.00 3.68 3.81
StackExchange SPLiCeBM25 5.07 (.07) 3.88 (.06) 3.32 (.04) 2.89 (.05) 3.60 (.05) 3.69 (.05)
SPLiCeCont 5.09 (.05) 3.91 (.03) 3.35 (.01) 2.93 (.01) 3.63 (.02) 3.73 (.01)
Baseline 5.14 3.94 3.36 2.94 3.65 3.74
Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. We perform three trainings using different training datesets and report the mean and standard deviation.
Long Context Method arXiv Code Code &
Data Python All arXiv
C SPLiCeBM25 5.463 $\pm$ .002 2.810 $\pm$ .002 2.942 $\pm$ .005 3.100 $\pm$ .004
SPLiCeCont 5.477 $\pm$ .005 2.824 $\pm$.001 2.957 $\pm$ .006 3.115 $\pm$ .006
SPLiCeRepo 5.474 $\pm$ .007 2.827 $\pm$ .006 2.958 $\pm$ .009 3.115 $\pm$ .009
Baseline 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
[…]
RoPe scale Training Method arXiv Code Code &
method data Haskell Python CUDA All arXiv
Naive C# SPLiCeBM25 6.25 (.08) 4.84 (.19) 3.55 (.11) 2.84 (.12) 3.72 (.13) 3.88 (.12)
SPLiCeRepo 6.25 (.08) 4.87 (.16) 3.56 (.10) 2.85 (.11) 3.74 (.11) 3.89 (.11)
Baseline 6.33 5.03 3.66 2.96 3.85 4.00
CodeLlama C# SPLiCeBM25 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.53 (.05) 3.34 (.06) 3.49 (.06)
SPLiCeRepo 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.54 (.04) 3.35 (.05) 3.50 (.05)
Baseline 5.76 4.37 3.27 2.58 3.40 3.55
YaRN C# SPLiCeBM25 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.55 (.06) 3.37 (.07) 3.52 (.06)
SPLiCeRepo 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.56 (.05) 3.38 (.06) 3.53 (.05)
Baseline 5.79 4.42 3.29 2.61 3.44 3.58
SPLiCe works with various context extension results Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension. FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. This allows for computationally efficient long-context fine-tuning. To show that our method works more generally, we have also evaluated it using more standard approaches. That is, we checked the effectiveness of the method when context extension is done in all layers, and parameters of Rotary Positional Encodings are either adjusted as in CodeLlama (Rozière et al., 2023), adjusted with YaRN (Peng et al., 2023), or left without changes. Table 3 shows the results.
Training is stable. We prepared three subsets of the C code dataset and ran the tested methods on each of them. In Table 3, we report the mean perplexity and its standard deviation. We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
3.3 Design Choices Analysis
Choice of k As already mentioned in the previous section, the choice of the retrieval method (BM25 or Contriver) plays a minor role. In Appendix C.1, we study the role of $k$. We first observe that its choice has a minor impact on the results, with $k=1$ being the best choice. We leave for further work a more detailed analysis of this hyperparameter.
Document ordering We discovered that the ordering of the documents collected by the retrieval procedure might be a subtle issue. We consider three choices of ORDER function in Algorithm 1: identity (do not permute the documents), reverse (reverse the order of documents) and a random shuffle. In Appendix C.1, we present results for $270$M models showing negligible differences. However, we observed differences for large models, elaborated in Section 4.2.4, in which random shuffling is the best choice for the CodeLlama context extension method.
Corpus chunking Performing retrieval from the whole corpus might be cumbersome due to its scale. However, we find that chunking the corpus into smaller parts still brings improvements. In particular, for the experiments with English Wikipedia and BM25 presented in Table 3, we used a manageable size of $2$GB.
4 Large-Scale Models
In this section, we show that SPLiCe can improve the long context performance of large-scale language models. To this end, we use $3$B and $7$B parameter models and tasks that test in-context learning (Section 4.2.1), question answering (Section 4.2.2), and in-context information retrieval capabilities (Section 4.2.3). We provide perplexity results in Appendix E.
We also emphasize that those improvements occur during a relatively short, compared to pre-training, fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B models were tuned on $2$B tokens. All of the models were pre-trained on $1$T tokens.
4.1 Experimental Setup
For $3$B experiments, we continue training OpenLLaMA $3$B v2 (Geng & Liu, 2023; Geng, 2023) on a 50/50 mixture of RedPajama prepared in the standard way and C prepared using SPLiCeBM25. For $7$B ones, we continue the training of OpenLLaMA $7$B v2 on 75/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCeBM25. StackExchange is part of RedPajama (TogetherComputer, 2023), and C data come from StarCoder (Li et al., 2023b).
We train with $32$K context length using $5.4$B tokens for $3$B models and $2$B for $7$B models (recall that OpenLLaMA models have the $2$K context length). We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023).
Regarding the context extensions method we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test six models:
[…]
where Baseline denotes the standard data preparation method. Hyperparameter details are located in Appendix A.
4.2 Experimental Results
We present perplexity measurements in Appendix E. In particular, following (Anthropic, 2023), we examine perplexity improvements in the function of the token position in the document. Our findings reveal that SPLiCe models demonstrate superior perplexity across all distances. Notably, the improvement in perplexity is more pronounced for tokens positioned later in the document, indicating enhanced language modeling capabilities.
For the remainder of this section, we focus on testing the in-context learning, question answering and information retrieval abilities on our models using downstream tasks.
4.2.1 In-Context Learning
To evaluate the improvements of in-context learning abilities, we use TREC (Li & Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks. For TREC, to test the long context utilization, we vary the number of in-context examples in $\{190,380,780,1560\}$ (which correspond roughly to $\{4,8,16,32\}$K context length). We sample these examples multiple times and measure classification accuracy on test set. In Table 4, we observe that SPLiCe improves the average performance across different context lengths and model sizes. We follow a very similar protocol for DBpedia, see Table 5, confirming that SPLiCe improves the context utilization. In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the domination of SPLiCe.
In Table 4 and Table 5, we report mean improvement $\Delta$ along with $95\%$ bootstrap confidence intervals. See Appendix I for related histograms.
Table 4: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001). For TREC, we average across $50$ sets of in-context examples for $3$B models and $10$ sets for $7$B models. We use $\Delta$ [confidence interval] to denote mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I).
TREC
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K (780) 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ]
$8$K (380) 66.5 75.8 9.3 [8.1,10.8]{8.1,10.8]}8.1 , 10.8 ]
$4$K (190) 65.1 72.8 7.7 [6.5,9.0]{6.5,9.0]}6.5 , 9.0 ]
[…]
$16$K (780) 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ]
$8$K (380) 72.1 76.2 4.1 [2.9,5.7]{2.9,5.7]}2.9 , 5.7 ]
$4$K (190) 66.9 69.7 2.8 [1.5,4.2]{1.5,4.2]}1.5 , 4.2 ]
[…]
$16$K (780) 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ]
$8$K (380) 76.6 76.5 -0.1 [1.0,1.1]{-1.0,1.1]}- 1.0 , 1.1 ]
$4$K (190) 68.1 69.1 1.0 [0.4,2.9]{-0.4,2.9]}- 0.4 , 2.9 ]
Table 5: Average performance on DBpedia (Lehmann et al., 2015). We average results across $40$ sets of in-context examples, with $\Delta$ [confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence interval. Due to the size of the evaluation dataset for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
DBpedia
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K (190) 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
[…]
$16$K (190) 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
[…]
$16$K (190) 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
Table 6: Average performance on HotPotQA (Yang et al., 2018). We average results across $7$ sets of in-context examples, with with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals.
HotPotQA
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
4.2.2 Question Answering
We also show that SPLiCe models have improved question answering capabilities. To this end, we use Qasper (Dasigi et al., 2021; Shaham et al., 2022) and HotPotQA (Yang et al., 2018). Qasper contains questions regarding research papers provided as a part of the input context. In Table 7, we observe that SPLiCe improves the two-shot performance. Similarly, we observe improvements on HotPotQA, see Table 6. We stress that in both cases, the results measure in-context capabilities as neither of the datasets was used during training.
Table 7: Two-shot F1 performance on the validation subset of Qasper (Dasigi et al., 2021). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
QASPER
Model
Context length
(#examples)
Baseline SPLiCe
[…]
4.2.3 Key Retrieval Performance
The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models, see Appendix D for the task details.
In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method. We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions. We provide the results for FoT models in Appendix D.
Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions. The details about this task can be found in Appendix D. Each position was evaluated using $500$ examples.
4.2.4 Shuffling
In our experiments with large $7$B CodeLlama models, we have noted slight improvements in handling long inputs when shuffling documents inside the context, that is, setting Order in the Algorithm 1 to be a random shuffle. We provide a comparison on TREC in Table 17.
5 Related Work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022). A recent work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, (Han et al., 2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of this approach in large-scale language modeling, specifically with OpenLLaMA $3$B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code. The gradient-based method presented in (Han et al., 2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
Concurrently to our research, (Shi et al., 2023b) showed a similar method for preparing the training data. However, in our work, we focus on tuning the models for long context, up to $32$K tokens, whereas (Shi et al., 2023b) focuses on training the models from scratch, with context length up to $8$K tokens. In addition to that, we provide ablations regarding context extension methods.
6 Limitations and Future Work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method can viewed as a general framework for organizing the documents into training examples. This opens multiple further research avenues.
Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work. Note that as BM25 is a syntactic retriever, it is not guaranteed that the documents returned by it are semantically related. Moreover, we have not tuned Contriever-MSMARCO, and it was provided only with a $512$ token long prefix of a document.
In order to prepare the training data, SPLiCe uses each document exactly once (we mask out each used document for further retrieval). However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces, such as paragraphs or sentences.
Scaling Our method requires further studies to understand its scaling properties beyond the range presented in the paper.
Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe. However, we leave more detailed studies of the design choices of SPLiCe with other context extension methods to future work.
Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a). This is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, answers, etc.) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
Data curation Using highly correlated samples has the potential to result in training instability. However, we noted no performance degradation during our experiments. We leave the study of how SPLiCe integrates with different data types for the future.
7 Conclusions
In this work, we present SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves the perplexity of the language modeling task and performance on downstream tasks. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
8 Impact Statement
This paper presents work whose goal is to improve the long context utilization ability of language models through structuring training data. It bears standard risks associated with advancing the field of LLMs (such as misuse and generation of malicious/misleading content). For a detailed overview, we refer to (Bender et al., 2021; Weidinger et al., 2021).
Acknowledgments
We gratefully acknowledge that our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, CI TASK) for providing computer facilities and support within computational grant numbers PLG/2023/016265. Piotr Miłoś was supported by National Science Center Poland under the grant agreement 2019/35/O/ST6/03464.
Appendix A Architecture
The architecture of our models is based on LLaMA (Touvron et al., 2023), and the architectural details can be found in Table 8. Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al., 2017) with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding (Su et al., 2021). For context extension, we use Focused Transformer (FoT)(Tworkowski et al., 2023), CodeLlama (Rozière et al., 2023) (CL) and YaRN (Peng et al., 2023). Table 9 presents the details about both standard and long-context pretraining.
Table 8: Architecture details. Focused Transformer context extension is applied in continued pretraining for $32$K context and evaluation.
Parameter/Model Size 270M 3B 7B
Vocab Size 32000 32000 32000
Embedding Size 1024 3200 4096
Num Attention Layers 12 26 32
Num Attention Heads 8 32 32
Head Size 128 100 128
MLP Hidden Size 4096 8640 11008
FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24]
Table 9: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B parameter OpenLLaMAv2 model (Geng, 2023). Subscript denotes that parameter was specific for a context extension method with FoT referring to (Tworkowski et al., 2023) and no-FoT to other methods (Naive, YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023)).
Stage Parameter/Model Size 270M 3B 7B
Pretraining Context 2K 2K 2K
Tokens 6.3B 1T 1T
Long Context Pretraining Context 32KFoT, 16Kno-FoT 32K 32K
Batch Size 128FoT, 16no-FoT 128 128FoT, 32no-FoT
Start Learning Rate 5e-5 1.5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6 1.5e-6
Warmup Steps 250 1000 1000FoT, 200no-FoT
Tokens 1B 5.4B 2B
Appendix B Detailed Results For Medium-Size Models
In this section, we extend results presented in Section 3. The training methodology is described in Section 3.1. Details about the number of tokens used for evaluation can be found in Table 18.
Tables 10 and 11 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 10 contains detailed results from training on C# and Python. Table 11 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix G). Both tables show that SPLiCe outperforms the Baseline by a significant margin.
The main advantage of SPLiCeBM25/SPLiCeCont over the SPLiCeRepo approach is that it can be used also for non-structured data. Table 12 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Table 13 considers models trained with YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023) and Naive (no adjustment to RoPE) context extension methods. Table 14 shows that a simple artificial extension of example length via random concatenation of documents does not help.
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways. For details about training please refer to Section 3.1.
Altered Train Data: C#
Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
ArXiv 5.52 5.53 5.65 5.53
C 2.39 2.40 2.50 2.40
C++ 2.60 2.61 2.74 2.62
CUDA 2.46 2.48 2.65 2.49
C# 1.82 1.82 1.90 1.82
Common Lisp 3.41 3.44 3.72 3.42
Dart 2.14 2.16 2.31 2.16
Emacs Lisp 8.41 8.46 8.85 8.41
Erlang 2.80 2.80 2.95 2.81
Fortran 3.68 3.71 4.05 3.72
Go 1.94 1.95 2.06 1.96
Groovy 3.01 3.03 3.25 3.04
Haskell 3.33 3.35 3.58 3.35
Java 2.09 2.10 2.22 2.10
Pascal 3.61 3.62 3.80 3.60
Python 2.90 2.91 3.07 2.91
Mean 3.26 3.27 3.46 3.27
Altered Train Data: Python
Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
ArXiv 5.47 5.49 5.57 5.48
C 2.38 2.39 2.45 2.39
C++ 2.61 2.63 2.71 2.63
CUDA 2.41 2.43 2.56 2.44
C# 1.98 1.99 2.06 2.00
Common Lisp 3.23 3.28 3.47 3.27
Dart 2.15 2.17 2.27 2.17
Emacs Lisp 7.94 7.98 8.28 7.93
Erlang 2.70 2.71 2.82 2.71
Fortran 3.42 3.46 3.72 3.46
Go 1.95 1.96 2.03 1.96
Groovy 2.97 2.99 3.13 2.99
Haskell 3.25 3.28 3.46 3.27
Java 2.12 2.12 2.20 2.13
Pascal 3.57 3.58 3.73 3.58
Python 2.53 2.53 2.62 2.54
Mean 3.17 3.19 3.32 3.18
Table 11: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix G and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
Altered Train Data: C
Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
Common Lisp 2.994 $\pm$ 0.072 3.042 $\pm$ 0.091 3.195 $\pm$ 0.078 3.043 $\pm$ 0.102
Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
Emacs Lisp 7.857 $\pm$ 0.017 7.851 $\pm$ 0.020 8.098 $\pm$ 0.027 7.840 $\pm$ 0.019
Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
Table 12: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods. For details about training, please refer to Section 3.1. Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
Altered Train Data: StackExchange
Eval/Method SPLiCeBM25 SPLiCeCont Baseline
ArXiv 5.07 5.09 5.14
C 2.68 2.69 2.70
C++ 3.02 3.04 3.06
CUDA 2.89 2.93 2.94
C# 2.27 2.28 2.29
Common Lisp 4.02 4.06 4.08
Dart 2.58 2.60 2.61
Emacs Lisp 9.55 9.67 9.69
Erlang 3.13 3.16 3.18
Fortran 4.28 4.34 4.38
Go 2.24 2.25 2.27
Groovy 3.62 3.66 3.68
Haskell 3.88 3.91 3.94
Java 2.43 2.45 2.45
Pascal 4.08 4.11 4.14
Python 3.32 3.35 3.36
Mean 3.69 3.73 3.74
Altered Train Data: Wikipedia
Eval/Method SPLiCeBM25 SPLiCeCont Baseline
ArXiv 5.64 5.65 5.73
C 2.65 2.67 2.71
C++ 2.98 3.01 3.07
CUDA 2.87 2.92 3.00
C# 2.22 2.24 2.29
Common Lisp 3.87 3.96 4.08
Dart 2.51 2.55 2.61
Emacs Lisp 9.38 9.45 9.63
Erlang 3.13 3.16 3.23
Fortran 4.23 4.32 4.49
Go 2.18 2.21 2.26
Groovy 3.49 3.55 3.67
Haskell 3.82 3.87 3.97
Java 2.39 2.41 2.46
Pascal 4.32 4.23 4.40
Python 3.26 3.30 3.37
Mean 3.68 3.72 3.81
Table 13: Perplexity results comparing different ways of organizing the same data for non-FoT models. All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C# code is organized in one of three ways.
Altered Train Data: C#
Context $16$K: CodeLlama
Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
ArXiv 5.74 5.76 5.74
C 2.66 2.70 2.66
C++ 2.79 2.83 2.79
CUDA 2.53 2.58 2.54
C# 1.91 1.93 1.91
Common Lisp 3.78 3.85 3.79
Dart 2.28 2.33 2.28
Emacs Lisp 8.29 8.41 8.30
Erlang 3.57 3.64 3.58
Fortran 3.93 4.01 3.95
Go 1.99 2.03 2.00
Groovy 2.95 3.01 2.96
Haskell 4.28 4.37 4.28
Java 2.31 2.35 2.31
Pascal 3.67 3.72 3.67
Python 3.22 3.27 3.22
Mean 3.49 3.55 3.50
Context $16$K: YaRN
Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
ArXiv 5.77 5.79 5.77
C 2.68 2.72 2.68
C++ 2.81 2.85 2.81
CUDA 2.55 2.61 2.56
C# 1.92 1.94 1.92
Common Lisp 3.83 3.92 3.84
Dart 2.30 2.36 2.30
Emacs Lisp 8.37 8.52 8.38
Erlang 3.60 3.67 3.61
Fortran 3.97 4.06 3.99
Go 2.00 2.04 2.01
Groovy 2.98 3.03 2.98
Haskell 4.32 4.42 4.32
Java 2.33 2.37 2.33
Pascal 3.69 3.75 3.69
Python 3.24 3.29 3.24
Mean 3.52 3.58 3.53
Context $16$K: Naive
Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
ArXiv 6.25 6.33 6.25
C 2.88 2.96 2.89
C++ 3.04 3.13 3.05
CUDA 2.84 2.96 2.85
C# 2.04 2.08 2.04
Common Lisp 4.40 4.56 4.39
Dart 2.50 2.60 2.51
Emacs Lisp 9.25 9.46 9.25
Erlang 3.98 4.10 4.00
Fortran 4.56 4.79 4.59
Go 2.14 2.21 2.16
Groovy 3.27 3.39 3.28
Haskell 4.84 5.03 4.87
Java 2.52 2.60 2.53
Pascal 4.05 4.20 4.10
Python 3.55 3.66 3.56
Mean 3.88 4.00 3.89
Table 14: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways. For details about training please refer to Section 3.1.
Altered Train Data: C
Eval/Method SPLiCeBM25 Baseline Random
ArXiv 5.46 5.55 5.55
C 2.13 2.17 2.18
C++ 2.40 2.47 2.47
CUDA 2.22 2.33 2.33
C# 1.94 2.02 2.02
Common Lisp 2.99 3.20 3.18
Dart 2.14 2.27 2.27
Emacs Lisp 7.86 8.10 8.09
Erlang 2.67 2.77 2.77
Fortran 3.31 3.55 3.56
Go 1.91 2.00 2.00
Groovy 3.01 3.15 3.15
Haskell 3.20 3.37 3.37
Java 2.07 2.16 2.16
Pascal 3.49 3.62 3.64
Python 2.81 2.93 2.93
Mean 3.10 3.23 3.23
Appendix C Ablations
C.1 SPLiCe Parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 15 indicates that $k=1$ is the best choice, though the differences are rather small. We found that changing the order of documents in training examples hardly matters for $270$M models. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
Table 15: Ablation of SPLiCe hyper-parameters. For each ablation, we have trained the same $270M$ parameter model using different data organization methods. Top-$k$ corresponds to the number of descendants chosen in the $\texttt{RETRIEVE}(d,k)$ step of the Algorithm 1. Reverse and shuffle correspond to the final order of examples $C$ returned by the Algorithm 1 (reverse – the order of documents in $C$ is reversed, shuffle – examples are shuffled.
Code
Method Top-$k$ Order C++ Haskell Python CUDA All
SPLiCeBM25 top 1 standard 2.38 3.20 2.82 2.23 2.93
reverse 2.38 3.21 2.82 2.23 2.93
top 2 standard 2.39 3.23 2.84 2.24 2.95
reverse 2.39 3.24 2.84 2.24 2.95
top 3 standard 2.39 3.24 2.84 2.25 2.97
reverse 2.39 3.25 2.84 2.25 2.96
shuffle 2.40 3.24 2.84 2.25 2.96
We also evaluated the influence of BOS and EOS tokens on the performance of trained models. As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively slightly decrease the number of separating tokens compared to the Baseline. However, in Appendix C.2 we included experiments showing that this has no impact on model performance.
C.2 Importance of Separating Tokens
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training examples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement). The results are shown in Table 16. We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains.
Table 16: Perplexity evaluation of two methods of organizing the data. Baseline – document equals training example. Random – concatenate documents into examples of length bounded by $120$K characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in Random does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation.
Long Context Method arXiv Code Code &
Data Python All arXiv
C Random 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
Appendix D Key-Value Retrieval Task
Figure 2 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in (Liu et al., 2023). This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
Extract the value corresponding to the specified key in the JSON object below. JSON data: {"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",  "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",  "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",  "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",  "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",  "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",  "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",  "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}  "4701aa05-c523-4b89-9700-64ab9c37c537":
We noted that FoT-trained models struggle with this task. This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits. Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 3, and 3. For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 3.
Figure 3: Performance on a smaller version of key-value retrieval task from (Liu et al., 2023). We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama (Rozière et al., 2023) method of context extension. Each position was evaluated using $500$ examples.
Appendix E Perplexity Improvements
In Figure 4 we present perplexity improvements of $3$B FoT SPLiCeBM25 over Baseline. Figure 5 shows the evolution of SPLiCe model perplexity during the training. We follow (Anthropic, 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
Figure 4: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps). We bucket tokens by their positions in the document and calculate the average. Each dot is the difference of the averages of the SPLiCe and Baseline models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 5: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 4 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
Appendix F Shuffling
Table 17: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001). We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was an identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
TREC
Model
Context length
(# of examples)
SPLiCe-no-shuf SPLiCe-shuf
$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
Appendix G Data Preparation
G.1 Evaluation Data
We have taken a random subset of arXiv from Proof-pile. For StarCoder data, we have downloaded up to $64$GB of each of the mentioned language subsets and performed a random 85/15 split for languages that we train on.
When evaluating the perplexity of the model, we skip documents that are shorter than the model context and truncate documents that are longer than that. Table 18 shows the number of tokens over which the perplexity was calculated.
Table 18: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents having not less than $c$ tokens and extract the $c$ tokens prefix.
Eval/Method 16K 32K 64K
ArXiv 16M 16M 16M
C 16M 16M 16M
C++ 16M 16M 16M
CUDA 16M 14M 8M
C# 16M 16M 16M
Common Lisp 16M 16M 16M
Dart 16M 16M 16M
Emacs Lisp 16M 15M 9M
Erlang 16M 16M 12M
Fortran 16M 16M 16M
Go 16M 16M 16M
Groovy 11M 4M 2M
Haskell 16M 16M 15M
Java 16M 16M 16M
Pascal 16M 16M 16M
Python 16M 16M 16M
G.2 Train Data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section G.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$K files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used to organize the data. For C, we have performed a run on three packs and provided results and standard deviation in Table 3. For large models, we run the methods on several packs and concatenate the results into a single dataset. For natural language datasets, we extract a random subset of documents.
Appendix H Faiss Parameters
Our experiments with SPLiCeCont utilize Faiss (Johnson et al., 2019) for fast approximate inner-product search. To be more precise, we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
Appendix I Detailed Accuracy Improvements
The performance of in-context learning depends much on the choice of the in-context examples. To study this in more detail we study the following random variable
$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{Baseline}}(c),$
where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{Baseline}}(c)$ are the accuracies of the model trained with SPLiCe and Baseline respectively on the random choice of in-context examples $c$. Below we report the histograms of $\delta(c)$. In Table 4 and Table 5 we report mean $\Delta$ and $95\%$ confidence intervals as $\Delta$ [confidence interval].
Figure 6 shows details about accuracy improvements on TREC when considering different numbers of in-context examples.
(a) Examples: $190$, Context $4$K
(b) Examples: $380$, Context $8$K
(c) Examples: $780$, Context $16$K
(d) Examples: $1560$, Context $32$K
Figure 6: Histograms of accuracy improvement of SPLiCeBM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the Baseline.
(a) Examples: $190$, Context $16$K, Model: $3$B FoT
(b) Examples: $380$, Context $32$K, Model: $3$B FoT
(c) Examples: $190$, Context $16$K, Model: $7$B FoT
(d) Examples: $380$, Context $32$K, Model: $7$B FoT
(e) Examples: $190$, Context $16$K, Model: $7$B CL
(f) Examples: $380$, Context $32$K, Model: $7$B CL
Figure 7: Histograms of accuracy improvement of SPLiCeBM25 over Baseline on DBPedia. We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
Appendix J Longer Context
In Table 19, we present the perplexity results for $270$M parameter models trained as described in Section 3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
[…]
Long Context Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
C# SPLiCeBM25 4.86 (.15) 2.60 (.19) 2.66 (.16) 2.32 (.19) 2.75 (.19) 2.88 (.19)
SPLiCeCont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
SPLiCeRepo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
Baseline 5.01 2.79 2.82 2.51 2.94 3.07
Appendix K Data Distribution Property
Chan et al. (2022); Han et al. (2023) shows that the ”burstiness” property of training data benefits a model’s generalisation. Specifically, ”burstiness” presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency. We also observe a burstiness property of data in SPLiCe, as shown in Table 20.
Table 20: Zipf’s coefficient of token frequency on Baseline and SPLiCe. A lower Zipf’s coefficient represents a more significant burstiness property.
Training Data Context Length Method Zipf’s Coefficient
C 2K SPLiCeBM25 ${1.667}_{(0.139)}$
Baseline ${1.717}_{(0.152)}$
StackEx 2K SPLiCeBM25 ${1.836}_{(0.088)}$
Baseline ${1.865}_{(0.074)}$
C 32K SPLiCeBM25 ${1.512}_{(0.055)}$
Baseline ${1.593}_{(0.025)}$
StackEx 32K SPLiCeBM25 ${1.643}_{(0.026)}$
Baseline ${1.664}_{(0.013)}$